### v1

# Fast Dynamic 3D Object Generation from a Single-view Video

Abstract
Generating dynamic three-dimensional (3D) object from a single-view video is challenging due to the lack of 4D labeled data. Existing methods extend text-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, but they are slow and expensive to scale (e.g., 150 minutes per object) due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this limitation, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly train a novel 4D Gaussian splatting model with explicit point cloud geometry, enabling real-time rendering under continuous camera trajectories. Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the same level of innovative view synthesis quality. For example, Efficient4D takes only 14 minutes to model a dynamic object.
11footnotetext:  Corresponding to lizhangfd@fudan.edu.cn
1 Introduction
Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time. Despite the significant advancement of three-dimensional (3D) object generation, existing works [28, 17, 4, 41] mostly consider static scenes or objects. With the availability of large-scale 3D datasets [6, 5], training generalizable models capable of directly generating multi-view images becomes possible [20, 21, 22]. These generated images can be turned into a 3D object through reconstruction techniques [25, 39]. By further augmenting these generated static objects with predefined animations [1], dynamic 3D content can be generated. However, this is highly limited due to the need for fine-grained meshes as well as rigid restrictions.
Refer to caption Refer to caption
Refer to caption Refer to caption
Input Generated dynamic objects
Figure 1: Examples of video-to-4D generation. Input: A brief video of a dynamic object, as represented by 3 frames per case; Output: Generated novel views at different timestamps.
Directly generating 4D object/scene content from text description has been recently attempted [36]. To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [25] via score distillation sampling [28] and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work [10] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as [36] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable in practice.
To address identified limitations, we formulate an efficient video-to-4D object generation method in a two-staged pipeline called Efficient4D. In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing filter into a previous multi-view image generator (e.g., SyncDreamer [21]) in tandem with frame interpolation. In the second stage, we use these training data to optimize a novel 4D Gaussian splatting model. This is an extension of the recent 3D Gaussian splatting [12], originally designed for static 3D scene representation, with the temporal dimension introduced additionally on top. Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs. To tackle the discontinuity between generated frames, we design a inconsistency-aware loss function. A couple of video-to-4D results are presented in Fig. 1.
Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem. (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches. (iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Guardian splatting model with the real-time rendering advantage.
(iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the same level of new view synthesis. Additionally, our method can also work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
2 Related work
4D representation
Efforts to synthesize videos with free viewpoint control in dynamic scenes have a well-documented history [46]. For example, pre-NeRF [25] approaches faced challenges in reconstructing intricate scene details. Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [29], which decouples geometry and motion, utilizing a canonical space and a learned deformation field. HyperNeRF [27] extends the deformation field to a higher dimension, employing the level-set method for smoother optimization. DeVRF [18] reconstructs the canonical space from multi-view dense observations in the initial frame and learns the deformation field with sparse views in subsequent timesteps, utilizing a voxel grid for representation.
Similarly, deformable 3D Gaussians [44] integrates the deformation field with 3D Gaussian Splatting for the scene’s geometric structure. DyNeRF [14] extends NeRF with an additional latent code to capture detailed temporal variations, while DynIBaR [16] extends the image-based rendering paradigm for representing long videos with complex camera and object motions. Another group of method [7, 3, 33] adopts tensor decomposition of 4D volumes to represent dynamic 3D scenes.
However, these representations often involve complex training schedules or suffer from slow convergence. In this work, drawing inspiration from 3D Gaussian Splatting [12], we propose an efficient 4D scene primitive for the representation of dynamic scenes.
3D generation
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion [28], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [17, 24, 32, 34, 4, 15, 38, 8, 42, 41, 45] and image-to-3D tasks [43, 37, 23, 30] across various dimensions.
However, SDS-based approaches often suffer from extended optimization times. Conversely, some efficient methods [26, 11, 20, 19, 21] have emerged. Notably, Point-E [26] and Shap-E [11] train models to directly generate 3D point clouds or meshes. Zero123 [20] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [19]. Importantly, SyncDreamer [21] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
4D generation
There are a few recent works dedicated for more challenging 4D object generation. For instance, MAV3D [36] deals with a text-to-4D problem by training Hexplane [3] with a video diffusion model and SDS loss. Instead of text input, Consistent4D [10] conditions the generation of 4D object over time on a monocular video with richer and more specific information. However, this work is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design. To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the model training speed by 10$\times$ whilst maintaining the quality of novel view synthesis.
Figure 2: Overview of our Efficient4D approach. Given as the input a brief video depicting a dynamic object in limited perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time. Efficient4D comprises two components: (A) Image sequence synthesis across views and timestamps, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images. (B) A novel 4D Gaussian representation model (d) that represents the scene with a number of Gaussian points. It can be trained efficiently and robustly under the confidence-aware (e) supervision on the generated image matrix.
3 Method
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video . The model, illustrated in Fig. 2, comprises two key components:
  • An image synthesis pipeline (Sec. 3.1) generates images across views and timestamps, ensuring sufficient geometry and temporal consistency.
  • A novel 4D Gaussian representation model (Sec. 3.2) efficiently utilizes the synthetic images for accurate dynamic object reconstruction and novel view synthesis.
3.1 Image synthesis across views and timestamps
Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Fig. 2(A)). Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency. Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Fig. 2(c)). This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [35, 21], offering comprehensive information for modeling a dynamic object.
To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video (Fig. 2(a)) and proceed to generate the remaining views. Our task involves generating multi-view consistent images from a single image for each row. Existing image-to-3D methods, such as SyncDreamer [21], can be leveraged for this purpose. However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images. To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
Specifically, SyncDreamer generates $N$ multi-view images $\{\mathbf{x}_{0}^{(1)},\dots,\mathbf{x}_{0}^{(N)}\}$ of a static object using a synchronized $N$-view noise predictor $\{\epsilon_{\theta}^{(n)}|n=1,\dots,N\}$ that predicts synchronized noise for noisy multi-view images $\mathbf{x}_{t}^{1:N}$. The noise predictor is conditioned on information correlated with all views. Cross-view conditioning is achieved through a spatial feature volume $\mathcal{V}\in\mathbb{R}^{F\times V\times V\times V}$ unprojected by $\mathbf{x}_{t}^{1:N}$ to inject 3D-aware features into the shared noise predictor, ensuring geometrical consistency across views for static moments.
To impose temporal consistency, the information from $\mathcal{V}$ at different timestamps is aggregated using a time-synchronous spatial volume we design here (Fig. 2(b)). A smoothing filter layer is introduced into the spatial volumes of different frames/timestamps, incorporating a weight vector $\mathbf{w}=(w_{-k},\dots,w_{0},\dots,w_{k})$. At each denoising step, time-synchronized spatial volumes $\tilde{\mathcal{V}}_{i}$ for each input frame $i\in\{1,2,\dots,M\}$ are constructed as:
$\tilde{\mathcal{V}}_{i}=\sum_{j=-k}^{k}w_{i+j}\mathcal{V}_{i+j}.$ (1)
This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, enhancing temporal consistency.
To further improve temporal resolution, video frame interpolation (e.g., RIFE [9]) is applied after generating the image matrix $\mathcal{D}$ as described above. The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames. This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
Analysis on temporal synchronization
For a clearer insight into our temporal synchronization design, we undertake a simplified experiment involving two input frames, each with its feature volumes labeled as $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. The fusion process is carried out by combining them as follows:
$\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (2)
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Fig. 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches 0.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio between 0.25 and 0.4, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
Refer to caption Refer to caption
Input Feature fusion
Figure 3: Example analysis on temporal synchronization. In this illustration, we manipulate the fusion ratio ($w$) across a range from 0 to 0.5 for the spatial feature volumes of two input frames. The results indicate that a moderate ratio can achieve superior outcomes by balancing both temporal consistency and motion independence.
3.2 4D generation through reconstruction
Recall that we desire a 3D dynamic object, we should not be content with the discrete images. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we formulate a novel 4D Gaussian representation model (Fig. 2(B)), departing from previous slow-to-train 4D reconstruction models [7].
4D Gaussian representation model
Our model builds upon the 3D Gaussian splatting technique introduced in [12], originally designed for static scene representation. We extend it to address the complexities of dynamic scenes, enabling the real-time synthesis of high-fidelity views in 4D space. The model represents each Gaussian defined as:
$G(\mathbf{p}|\mathbf{\mu},\Sigma)=e^{-\frac{1}{2}\left(\mathbf{p}-\mathbf{\mu}\right)^{\top}\Sigma^{-1}\left(\mathbf{p}-\mathbf{\mu}\right)},$ (3)
where $\mathbf{\mu}\in\mathbb{R}^{4}$ is the mean vector, and $\Sigma\in\mathbb{R}^{4\times 4}$ is the anisotropic covariance matrix. The input $\mathbf{p}=(\mathbf{x},t)\in\mathbb{R}^{4}$ represents a spacetime position with a spatial coordinate $\mathbf{x}$ and time $t$. The covariance matrix $\Sigma$ decomposes into a diagonal scaling matrix $S\in\mathbb{R}^{4\times 4}$ and a rotation matrix $R\in\mathbb{R}^{4\times 4}$ through $\Sigma=RSS^{\top}R^{\top}$. The 4D rotation $R$ is represented by a pair of iso-rotations, each characterized by a quaternion.
In contrast to the 3D counterpart, our formulation incorporates the time dimension, allowing coherent modeling of dynamics in both space and time.
Rendering
For rendering, each Gaussian includes opacity $\alpha$ and view-dependent color $\mathbf{c}$ represented by spherical harmonics (SH). Given an arbitrary view $\mathcal{I}$ defined by intrinsic and extrinsic parameters, we render the pixel at position $(u,v)$ with timestamp $t$ by blending visible Gaussians:
$\mathcal{I}(u,v,t)=$ $\sum_{i}f_{i}(u,v,t)\alpha_{i}\mathbf{c}_{i}(d)T_{i}$ (4)
with $\text{with }T_{i}=$ $\prod^{i-1}_{j=1}(1-f_{j}(u,v,t)\alpha_{j}).$ (5)
Here, $i$ indexes the visible Gaussians sorted by depth, $d$ refers to the direction of the pixel under the view $\mathcal{I}$, and $f_{i}(u,v,t)$ denotes the influence of a Gaussian on this position. To obtain the influence, unlike 3D Gaussian Splatting [12] which directly projects 3D Gaussians to image space, we need condition 4D Gaussians on time and then project them. More specifically, $f(u,v,t)$ is expressed by:
$f(u,v,t)=G_{t}(t|\mathbf{\mu}_{t},\Sigma_{t})G_{{P}(\mathbf{x})|t}(u,v|{P}(\mathbf{\mu}_{\mathbf{x}|t}),{P}(\Sigma_{\mathbf{x}|t})),$ (6)
where $G_{t}$ is the marginal distribution of the 4D Gaussian in time, and $G_{\text{P}(\mathbf{x})|t}$ is the projected version of the conditional 3D Gaussian with
$\mathbf{\mu}_{\mathbf{x}|t}$ $=\mathbf{\mu}_{\mathbf{x}}+\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}(t-\mathbf{\mu}_{t})$ (7)
$\Sigma_{\mathbf{x}|t}$ $=\Sigma_{\mathbf{x}}-\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}\Sigma_{t,\mathbf{x}}.$ (8)
The projection operation $P$ [12, 47] projects the world point $\mathbf{\mu}_{\mathbf{x}|t}$ to image space and transforms the covariance by ${P}(\Sigma_{\mathbf{x}|t})=JW\Sigma_{\mathbf{x}|t}W^{\top}J^{\top}$ where $W$ is the extrinsic matrix of $\mathcal{I}$ and $J$ is the Jacobian of the affine approximation of the projective transformation.
Optimization
During model training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original objective function, as presented in [12], assumes clean training data, which may not be applicable to synthetic data with inherent imperfections. To address this, we introduce a confidence-aware objective loss formulation defined as:
$\mathcal{L}=\lambda_{\text{rgb}}\mathcal{C}_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{C}_{\text{ssim}}\mathcal{L}_{\text{ssim}},$ (9)
where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [40], $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter, and $\mathcal{C}_{\text{rgb/ssim}}$ is the confidence score of a generated image $\mathbf{I}$ calculated as
$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}})$ (10)
where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by the optical flow method. In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions.
4 Experiments
Input samples
Refer to caption Refer to caption Refer to caption Refer to caption
Consistent4D [10] ~165 min
Refer to caption Refer to caption Refer to caption Refer to caption
Ours images ~2 min
Refer to caption Refer to caption Refer to caption Refer to caption
Ours geometry ~12 min
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative comparisons. We show four images for each method, captured from two viewpoints at two timestamps.
4.1 Implementation details
Competitor
Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [10]. In comparison, we obtained the results of Consistent4D by referencing its original paper and project page since the source code has not been publicly released.
Evaluation videos
To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. Specifically, we focused on 20 sequences: alpaca, astronaut, bird, butterfly, cat, corgi, dinosaur, dog, dragon, duck, flower, guard, panda, penguin, rabbit, robot, snail, Spiderman, yellow face, yoxi.
For dragon, guard, and yoxi, 3D animated models were obtained from Sketchfab [2]. We rendered front-view videos and corresponding masks at 30 fps using Blender. The yellow face sequence was collected from the internet. To facilitate comparison with Consistent4D [10], we also conducted experiments on the videos used in their work [10].
Each sequence consists of monocular videos captured by a stationary camera, with durations ranging from approximately 1s to 6s. We uniformly selected 2-32 frames from each video as keyframes and segmented their foreground parts using SAM [13].
Table 1: Quantitative comparisons. The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering. ${}^{\dagger}$: Reported in [10] with one V100 GPU. ${}^{\ddagger}$: Measured on K-Planes [7], which partially serves as the fundamental 4D representation of Consistent4D [10].
Method CLIP-similarity $\uparrow$ Generation time $\downarrow$ FPS $\uparrow$ Explicit geometry
Consistent4D [10] 0.8438 165 ${}^{\dagger}$${}^{\ddagger}$
Efficient4D 0.8601 14 min ~1000
Implementation
For time-synchronous spatial volume, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$. In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian. Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 30,000-50,000. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 9,000 iterations with a batch size of 4. Gaussians are densified every 50 iterations, with densification halted at the 7,000th iteration. All other hyperparameters remain consistent with those in [12]. For speed efficiency, our proposed method only costs about 2 minutes for image generation and 12 minutes for reconstruction on one A6000 GPU.
4.2 Qualitative evaluation
We present qualitative comparisons in Fig. 4. In geometry rendering, we extract points at the mean of Gaussians. We also eliminate white background points and transparent points with low opacity. When assessing texture quality, it is important to note that both methods fall under the category of lifting 2D to 4D. However, Consistent4D produces watercolor-like images with low fidelity, which can be attributed to conflicts from multiple supervisory signals during prolonged optimization. In contrast, our method excels in directly generating high-quality 2D images. Subsequently, our method’s reconstruction stage yields explicit point clouds and exhibits less degradation in rendering quality compared to the SDS-based Consistent4D. Overall, our method consistently outperforms Consistent4D in most cases.
4.3 Quantitative evaluation
In Tab. 1, we compare Consistent4D and our method based on average generation time, rendering time, and CLIP-similarity scores [31] for input image sequences and rendered images. Measurements used the image sequences described in Sec. 4.1. Consistent4D’s CLIP-similarity evaluation is limited to RGB images from 2-3 selected views in their paper and project page, while our evaluation includes 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps, making our CLIP-similarity score more robust.
We draw several observation: (1) Our method achieves superior image quality. (2) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement. (3) Our efficient 4D Gaussian representation and splatting based rendering enable real-time rendering and provide explicit point cloud geometry.
More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization in fewer iterations, and (2) each iteration in our method requiring less time compared to Consistent4D. This time efficiency is further enhanced by avoiding recurrent evaluation in the diffusion model and prolonged forward rendering of the inefficient 4D representation as needed in the competitor at every iteration.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Input Generated images Geometry
Figure 5: Given only two input frames, our method is able to generate point clouds and temporal continuous images from novel views.
4.4 Sparse input evaluation
We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames. In contrast, Consistent4D fails to operate effectively under such conditions. In such cases, we set $w=0.25$ in Eq. 2. As illustrated in Fig. 5, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the video-to-4D task.
4.5 Ablation studies
Single-view
Refer to caption Refer to caption Refer to caption Refer to caption
No time-sync
Refer to caption Refer to caption Refer to caption Refer to caption
No interp
Refer to caption Refer to caption Refer to caption Refer to caption
Full setting
Refer to caption Refer to caption Refer to caption Refer to caption
Generated images Reconstructed Generated images Reconstructed
Figure 6: Ablation study on multi-view images generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
Generated images
K-Plane [7] ~2 hrs
No conf ~12 min
Full setting ~12 min
Reconstructed novel views
Figure 7: Ablation study on 4D scene representation and confidence maps. The images follow a chronological order from left to right.
We performed ablation studies to assess the influence of different components in our approach. We compared our full method against five baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded; (4) Confidence maps are excluded; (5) 4D Gaussian splatting is replaced with K-Plane [7]. The results corresponding to these settings can be found in Fig. 6 and Fig. 7.
Importance of synthetic data.
We compare our generated image matrix with utilizing only input video using the same proposed 4D Gaussian representation model. As shown in the first row of Fig. 6, when only relying on a single-view video, the model cannot produce any meaningful results for novel views. This indicates the importance of constructing proper training data.
Effect of time-synchronous spatial volume.
We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and fourth rows in Fig. 6. Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry. In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation.
Effect of frame interpolation.
As illustrated in the contrast between the third and fourth rows in Fig. 6, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views. This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
Effect of confidence-aware in objective.
The integration of a confidence-aware loss in our design serves as a strategy to mitigate training data noise. This is evident in the contrast between the third and fourth rows in Fig. 7. The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality.
Integrating existing 4D representation model.
Our synthetic training data is versatile and supports the optimization of 4D representation models, such as K-Plane [7]. In the comparison shown in Fig. 7, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes. Notably, our method surpasses K-Plane in speed, achieving a 10$\times$ improvement. Additionally, as emphasized in Sec. 4.3, our model enables real-time rendering with explicit point cloud geometry.
5 Conclusion
This study introduces a new framework, Efficient4D, designed for generating efficiently dynamic 4D objects seamlessly from monocular videos captured by a stationary camera. The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions using an innovative 4D Gaussian splatting model. Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 10 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results. Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.
Limitations.
Our local smoothing approach in sliding window style with our image sequence generator poses challenges in handling long-duration videos. To address this issue, replacing the smoothing filtering layer with a learnable attention layer featuring global receptive fields is a viable solution, particularly when dealing with extensive 4D datasets. Additionally, handling long videos may necessitate significant GPU memory, which can be mitigated by employing multi-GPUs or CPUs, albeit with an increase in processing time.